OpenStudy (anonymous):

i have problem in calculating MRS (marginal rate of substitution) from MU(marginal utility). please explain with example ?

OpenStudy (anonymous):

Well, I hope this helps. MU of a good X1 measures the contribution of its one EXTRA unit to the Total Utility (U). MRS shows us HOW MUCH of the good X1 we have to gave up for the obtaining of the EXTRA unit of X2 good (SEE THE 2nd relation below). 1) MRS is equal to the (negative) relation of prices of two goods: \[MRS=-\frac{ p1 }{ p2 }\] 2) MRS is equal to the RELATIVE change in quantaty of good X1 and good X2: \[MRS= \frac{ DeltaX1 }{ DeltaX2 }\] 3) MRS is equal to the (negative) relation od MU of good X1 and MU of the X2 good: \[MRS= - \frac{ MU1 }{ MU2 }\] \[MU1=\frac{ deltaUtility }{ deltaX1 }\] - it is the same for MU2 and X2 When you sum this up. it is easy to understand and calculate MRS from MU. \[\frac{ DeltaX1 }{ DeltaX2 }= \frac{ p1 }{ p2 }=\frac{ MU1 }{ MU2}\]

OpenStudy (anonymous):